The weak acid HA is 4 % ionized (dissociated) in a 0.10 M solution.

a)What is Ka for this acid?

b) What is the pH of this solution?
Answer #1

concentration = 0.10 M

% ionization = 4 %

a)

% ionization = ([H+] / C ) x 100

4 = ([H+] / 0.10) x 100

[H+] = 4.0 x 10^-3 M

[H+] = sqrt (Ka x C)

4.0 x 10^-3 = sqrt (Ka x 0.10)

Ka = 1.6 x 10^-4

b)

pH = -log [H+]

    = -log (4 x 10^-3)

pH = 2.40